There are two numbers. When four times the larger number is added to three times the smaller number, the result is 26. If three times the larger number is decreased by twice the smaller, the result is 11. Find the numbers.
Answer by mananth(16946)

let the bigger number be x and smaller be y
Condition I
4x+3y=26...................(1)
condition II
3x-2y=11....................(2)
Eliminate y
multiply (1)by 2
Multiply (2) by 3
8 x 6 y = 52
9 x -6 y = 33
Add the two equations
17 x = 85
/ 17
x = 5
plug value of x in (1)
4 x + 3 y = 26
20 + 3 y = 26
3 y = 26 -20
3 y = 6
y = 2